(((A stamp collection consists of 3, 8 and 15 cent stamps. The number of 8 cent stamps is one less than triple the number of 3 cent stamps. The number of 15 cent stamps is six less that the number of 8 cent stamps. The total value of all the stamps is $2.47. Find the number of 8 cent stamps in the collection.)))
I have tried every way I can think of to make an equation to solve this problem. I have been unable to do so. If I could get some help to come up with an equation for this problem I would be most grateful.

Answer by stanbon(48558)

A stamp collection consists of 3, 8 and 15 cent stamps.
The number of 8 cent stamps is one less than triple the number of 3 cent stamps.
Eq:; e = 3t-1
---------------------
The number of 15 cent stamps is six less that the number of 8 cent stamps.
Eq: f = e-6
---------------------
The total value of all the stamps is $2.47.
Eq: 3t + 8e + 15f = 247 cents
-----------------------------------
Find the number of 8 cent stamps in the collection.
---
Rearrange the three equations and solve for "e":
3t - e + 0 = 1
0t + e - f = 6
3t + 8e+15f = 247
-----------------------
Subtract the 1st from the 3rd
3t - e + 0 = 1
0t + e - f = 6
0t + 9e+15f = 246
---------------------------
Add 15 times the 2nd to the 3rd
3t - e + 0 = 1
0t + e - f = 6
0t +24e+ 0f = 336
------------------
Solve the 3rd for "e":
24e = 336
e = 14 (# of 8 cent stamps)
==============================
Cheers,
Stan H.